OpenStudy (anonymous):

2sin^2 theta + 3cos theta -3 =0 tyler, answer meee!! LOL

OpenStudy (anonymous):

\[\sin^2 \theta = 1 - \cos^2 \theta\]

OpenStudy (anonymous):

\[2(1-\cos^2 \theta)- 3 \cos \theta -3 = 0\]

OpenStudy (anonymous):

\[2-2\cos^2 \theta - 3 \cos \theta -3 = 0\]

OpenStudy (anonymous):

\[-2 \cos^2 \theta - 3\cos \theta-1 = 0\]

OpenStudy (anonymous):

times everything by -1 \[2\cos^2\theta + 3\cos \theta +1 = 0\]

OpenStudy (anonymous):

factor \[(2 \cos \theta + 1 )( \cos \theta + 1 ) = 0\]

OpenStudy (anonymous):

solve the two parts... (2cosθ+1) = 0 or (cosθ+1)=0 cos θ =- .5 or cos θ = -1

OpenStudy (anonymous):

θ = 120, 240, 180

OpenStudy (anonymous):

-fin-

OpenStudy (anonymous):

hmm...

OpenStudy (anonymous):

is that the pythagorean id?

OpenStudy (anonymous):

\[\sin^2\theta=1−\cos^2\theta\] is Pythagorean's Identity

OpenStudy (anonymous):

you meant -1 not equals 1.. right?

OpenStudy (anonymous):

if you flip it around, \[1 = \sin^2\theta + \cos ^2\theta\]

OpenStudy (anonymous):

ooh.